### v1

# Pilot Contamination-Aware Graph Attention Network for Power Control in CFmMIMO

Abstract
Optimization-based power control algorithms are predominantly iterative with high computational complexity, making them impractical for real-time applications in cell-free massive multiple-input multiple-output (CFmMIMO) systems. Learning-based methods have emerged as a promising alternative, and among them, graph neural networks (GNNs) have demonstrated their excellent performance in solving power control problems. However, all existing GNN-based approaches assume ideal orthogonality among pilot sequences for user equipments (UEs), which is unrealistic given that the number of UEs exceeds the available orthogonal pilot sequences in CFmMIMO schemes. Moreover, most learning-based methods assume a fixed number of UEs, whereas the number of active UEs varies over time in practice. Additionally, supervised training necessitates costly computational resources for computing the target power control solutions for a large volume of training samples. To address these issues, we propose a graph attention network for downlink power control in CFmMIMO systems that operates in a self-supervised manner while effectively handling pilot contamination and adapting to a dynamic number of UEs. Experimental results show its effectiveness, even in comparison to the optimal accelerated projected gradient method as a baseline.
Index Terms:
Graph attention network, power control, pilot contamination, cell-free massive MIMO.
I Introduction
Cellular massive multiple-input multiple-output (MIMO), usually enabled by multi-antenna panel arrays, is the backbone of today’s 5G communication technology. However, a persistent problem with cellular massive MIMO deployments is the poor service quality experienced by the cell-edge user equipments (UEs). Cell-free massive MIMO (CFmMIMO) has been proposed as a promising architecture for next-generation wireless communications. CFmMIMO eliminates the cellular boundaries and instead deploys massive distributed access points (APs) across a large coverage area [1]. These APs cooperatively serve UEs within the coverage area without the constraints of cell boundaries, ensuring fairness in service quality across UEs at different locations and reducing interference through efficient beamforming enabled by the coordination of a massive number of distributed antennas. As a result, CFmMIMO can achieve a higher network spectral efficiency (SE) compared to traditional cellular schemes [2, 3].
To fully realize the benefits of CFmMIMO, several challenges must be addressed. In particular, power control [^1] We focus on downlink power control in this paper. and pilot allocation are critical. With an increasing number of UEs, especially in CFmMIMO systems, the limited number of orthogonal pilot sequences becomes insufficient, leading to pilot contamination. Pilot contamination significantly degrades the communication quality, and thus, its impact cannot be overlooked when addressing power control problems. Traditional power control algorithms are predominantly optimization-based, with computational complexity increasing sharply as the numbers of APs and UEs grow. For CFmMIMO systems with a large number of APs and UEs, existing algorithms lead to high computational costs and processing time, which runs counter to the demands for high efficiency and low latency.
To address this challenge, learning-based algorithms have been proposed to compute power allocation coefficients with significantly improved efficiency. Various neural network models, including convolutional neural networks [4, 5], transformer neural networks [6, 7], and graph neural networks (GNNs) [8, 9, 10, 11, 12] have been explored using supervised or self-supervised training approaches. GNNs have demonstrated excellent performance in solving power control problems thanks to their structural similarity to wireless communication networks [9, 10]. As mentioned earlier, pilot contamination cannot be ignored when addressing power control. However, to the best of our knowledge, all existing GNN-based methods assume the use of ideally orthogonal pilot sequences, which is a big stretch especially in CFmMIMO systems. To bridge this gap, we propose a pilot contamination-aware graph attention network to handle power control in CFmMIMO systems. The proposed network model utilizes a GNN enhanced with attention mechanism [6, 7] to achieve power control with the awareness of pilot contamination, and is able to handle dynamic scenarios with a varying number of UEs without requiring retraining. Besides, the network is trained in a self-supervised manner, reducing the workload of computing target power control coefficients. Experimental results are presented to compare the proposed method against the first-order accelerated projected gradient (APG) method described in [13] and transformer in [7].
II System Model
II-A Cell-Free Massive MIMO
Consider the downlink power control in a CFmMIMO system, where all UEs are served by APs using common time-frequency resources, and all APs are synchronized. The system operates in time-division duplex mode. The system consists of $M$ APs and $K$ single-antenna UEs randomly distributed over a large coverage area, with each AP equipped with $N$ antennas. Assuming that channels are block-fading and remain constant over a coherence block of $T_{\rm c}$ symbols, the uplink channel from the $k$th UE to the $m$th AP can be expressed as
$\mathbf{g}_{mk}=\sqrt{\beta_{mk}}{\mathbf{h}_{mk}},$ (1)
where $\beta_{mk}$ denotes the large-scale fading coefficient, and $\mathbf{h}_{mk}\in\mathbb{C}^{N}$ denotes the small-scale coefficient vector for an $N$-antenna array. The entries of $\mathbf{h}_{mk}$ are independent and identically distributed (i.i.d.) complex Gaussian random variables with zero mean and unit variance, i.e., $\mathcal{CN}(0,1)$.
[…]
In the uplink training phase, the received signal at the $m$th AP can be expressed as
$\mathbf{Y}_{{\rm up},m}=\sqrt{\zeta_{\rm p}T_{\rm p}}\sum_{i=1}^{K}\mathbf{g}_{mi}\bm{\psi}_{i}^{\rm H}+\mathbf{W}_{{\rm up},m},$ (2)
where $\zeta_{\rm p}$ denotes the normalized transmit signal-to-noise ratio per pilot symbol, $\left(\cdot\right)^{\rm H}$ represents the conjugate transpose operation, and $\mathbf{W}_{{\rm up},m}\in\mathbb{C}^{N\times T_{\rm p}}$ is the noise matrix whose entries follow i.i.d. $\mathcal{CN}(0,1)$. The minimum mean square error (MMSE) estimate of the channel between the $m$th AP and the $k$th UE can be derived as [13]
$\hat{\mathbf{g}}_{mk}=\frac{\sqrt{\zeta_{\rm p}T_{\rm p}}\beta_{mk}}{1+\zeta_{\rm p}T_{\rm p}\sum_{i=1}^{K}\beta_{mi}\left|\bm{\psi}_{i}^{\rm H}\bm{\psi}_{k}\right|^{2}}\mathbf{Y}_{{\rm up},m}\bm{\psi}_{k},$ (3)
[…]
$\overline{g}_{mk}=\mathbb{E}\left[\left|\hat{\mathbf{g}}_{mk}[n]\right|^{2}\right]=\frac{\zeta_{\rm p}T_{\rm p}\beta_{mk}^{2}}{1+\zeta_{\rm p}T_{\rm p}\sum_{i=1}^{K}\beta_{mi}\left|\bm{\psi}_{i}^{\rm H}\bm{\psi}_{k}\right|^{2}}.$ (4)
APs cooperate to deal with power control with the knowledge of large-scale fading information through the central processing unit, while each AP works independently to apply a beamforming scheme with the estimated small-scale fading coefficients. Denote the power control coefficient matrix as $\mathbf{M}\in\mathbb{R}^{M\times K}$, and the $(m,k)$-th matrix element, $\mu_{mk}$, is the power control coefficient for the signal transmitted from the $m$th AP to the $k$th UE. According to the practical constraint on the total transmit power per AP, the feasible set of power control coefficients can be expressed as
$\mathcal{S}=\left\{\mathbf{M}\mid\mu_{mk}\geq 0;\left\|\bm{\mu}_{m}\right\|^{2}\leq\frac{1}{N},\forall m,k\right\},$ (5)
where $\bm{\mu}_{m}$ is the $m$th row of $\mathbf{M}$. Denoting the data symbol transmitted to the $k$th UE as $s_{k}$ with $\mathbb{E}\left[\left|s_{k}\right|^{2}\right]=1$, and $\zeta_{\rm d}$ as the maximum downlink transmit power per symbol normalized to noise power at each AP, the downlink signal from the $m$th AP using conjugate beamforming can be expressed as
$\mathbf{x}_{m}=\sqrt{\zeta_{\rm d}}\sum_{k=1}^{K}\frac{\mu_{mk}}{\sqrt{\overline{g}_{mk}}}\hat{\mathbf{g}}_{mk}^{*}s_{k},$ (6)
where $\left(\cdot\right)^{*}$ denotes conjugate operation. Using the use-and-then-forget bounding technique, the signal-to-interference-plus-noise ratio for the $k$th UE can be derived as [13]
$\gamma_{k}=\frac{\zeta_{\rm d}\left(\overline{\bm{\mu}}_{k}^{\rm T}\bm{\nu}_{kk}\right)^{2}}{\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{K}\zeta_{\rm d}\left(\overline{\bm{\mu}}_{i}^{\rm T}\bm{\nu}_{ik}\right)^{2}+\frac{\zeta_{\rm d}}{N}\sum_{i=1}^{K}\left\|\overline{\mathbf{B}}_{k}\overline{\bm{\mu}}_{i}\right\|^{2}+\frac{1}{N^{2}}},$ (7)
where $\overline{\bm{\mu}}_{k}$ denotes the $k$th column of $\mathbf{M}$, $(\cdot)^{\rm T}$ is the transpose operator, $\overline{\mathbf{B}}_{k}$ is a diagonal matrix with diagonal elements $\sqrt{\beta_{1k}},\ldots,\sqrt{\beta_{Mk}}$, and the $m$th element of $\bm{\nu}_{ik}\in\mathbb{R}^{M}$ is
$\bm{\nu}_{ik}[m]=\left|\bm{\psi}_{i}^{\rm H}\bm{\psi}_{k}\right|\sqrt{\overline{g}_{mi}}\frac{\beta_{mk}}{\beta_{mi}}.$ (8)
The downlink network SE of the $k$th UE is then
$\mathrm{SE}_{k}=\left(1-\frac{T_{\rm p}}{T_{\rm c}}\right)\log_{2}\left(1+\gamma_{k}\right).$ (9)
II-B Optimization Problem Formulation
Define $\mathbf{B}\in\mathbb{R}_{+}^{M\times K}$ as the large-scale fading coefficient matrix with $\beta_{mk}$ as the $(m,k)$-th element, and let $\bm{\Phi}\in\mathbb{R}^{K\times K}$ denote the pilot allocation matrix, where the $(i,j)$-th element is given by $\left|\bm{\psi}_{i}^{\rm H}\bm{\psi}_{j}\right|$. Here $\mathbb{R}_{+}^{M\times K}$ represents the set of matrices of dimension $M\times K$ with non-negative real entries. From (7) and (9), SE depends on both the power control coefficients $\mathbf{M}$ and pilot allocation matrix $\bm{\Phi}$, given a large-scale fading matrix $\mathbf{B}$. When $T_{\rm p}\geq K$, pilot sequences are orthogonal across all UEs, making their impact on SE negligible. However, in massive MIMO scenarios, $T_{\rm p}<K$ often occurs. Therefore, incorporating the information of pilot assignment is essential for optimizing power control under pilot contamination.
Max-min fairness maximization is the objective given by
$\begin{array}[]{rrr}\underset{\mathbf{M}}{\operatorname{maximize}}\left\{\underset{1\leq k\leq K}{\operatorname{min}}\mathrm{SE}_{k}(\mathbf{M};\bm{\Phi},\mathbf{B})\right\}.\end{array}$ (10)
By adopting a smoothing technique with smoothing parameter $\lambda$, the objective function in (10) can be reformulated as [13]
$u(\mathbf{M};\bm{\Phi},\mathbf{B})=-\frac{1}{\lambda}\log\left(\frac{1}{K}\sum_{k=1}^{K}\exp\left({-\lambda\,\mathrm{SE}_{k}(\mathbf{M};\bm{\Phi},\mathbf{B})}\right)\right).$ (11)
Define a deep neural network as a map, $\mathbf{M}=f(\bm{\Phi},\mathbf{B};\mathbf{\Theta})$. Here $f$ maps $\bm{\Phi}$ and $\mathbf{B}$ to the power control matrix $\mathbf{M}$, and $\mathbf{\Theta}$ represents the trainable weights of the network. Our goal is to determine the optimal weights $\mathbf{\Theta}_{\rm opt}$ such that the neural network generates the power control coefficient matrix $\mathbf{M}$ to maximize the objective function $u$ in (11).
III Graph Attention Network for Power Control
III-A Graph Representation
Consider a heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ composed of a set of nodes $\mathcal{V}$ and edges $\mathcal{E}$. Each node $v$ represents an AP-UE pair and is mapped by $\varphi(m,k)=v$, where $\varphi$ is the mapping function from an AP-UE pair to its corresponding node, and $v=M(k-1)+m$. To enable a single power control network capable of handling a varying number of UEs, $K$ is set as the maximum number of UEs, $K_{\rm max}$, within the coverage area. If the actual number of UEs, $K_{\rm act}$, is smaller than $K_{\rm max}$, an extra $K_{\rm max}-K_{\rm act}$ UEs are added as padding. To differentiate between actual and padded UEs, each node is assigned a binary attribute, ${\rm Mask}$, and ${\rm Mask}_{i}=1$ indicates that the UE associated with the $i$th node is served, while ${\rm Mask}_{i}=0$ denotes a padded UE. The bidirectional edge between the $i$th and $j$th nodes is denoted as $e_{ij}$ for $i\neq j$. Edges are categorized based on whether the connected nodes share the same AP or UE, or neither:
  • AP-type edges ($e_{ij}^{\rm AP}$): Nodes share the same AP.
  • UE-type edges ($e_{ij}^{\rm UE}$): Nodes share the same UE.
  • No connection: If two nodes share neither an AP nor a UE, no edge exists between them.
To incorporate pilot contamination awareness, the pilot allocation matrix $\bm{\Phi}$ is included as an attribute for AP-type edges, capturing the orthogonality of allocated pilot sequences among UEs that share the same APs. Specifically, the $(i,j)$-th element of $\bm{\Phi}$, denoted as $\psi_{ij}$, is assigned to those AP-type edges whose two connected nodes correspond to the $i$th and $j$th UEs. Suppose there are $T$ attention layers in the graph structural network, and the input and output features of the $i$th node at the $t$th graph attention layer are represented by $\mathbf{x}_{i}(t-1)$ and $\mathbf{x}_{i}(t)$, respectively, where $t=1,\ldots,T$. The characteristics of the proposed graph structure, including the node representation $\mathbf{x}_{i}(t)$, the node attribute ${\rm Mask}_{i}$, the AP-type edges with pilot contamination information $\psi_{ij}$, and the UE-type edges, are illustrated in Fig. 1. This structure enables the network to effectively model the relationship between APs and UEs while accounting for pilot contamination.
Figure 1: Illustration of graph characteristics.
III-B The Proposed Network Model
1) Preprocessing Stage: To accommodate a varying number of UEs, the inputs $\mathbf{B}$ and $\bm{\Phi}$ are first padded. Specifically, zero-padding is applied to generate $\mathbf{B}_{\rm padded}$ of size $M\times K_{\rm max}$ and $\bm{\Phi}_{\rm padded}$ of size $K_{\rm max}\times K_{\rm max}$. For simplicity of notation, $\mathbf{B}$ and $\bm{\Phi}$ are used to refer to their padded counterparts, and $K=K_{\rm max}$ in the following discussion unless explicitly stated.
Given that the values of large-scale fading coefficients span different orders of magnitude, the natural logarithm operation is applied to each element of $\mathbf{B}$, resulting in $\mathbf{B}_{\rm log}$. Following this, $\mathbf{B}_{\rm log}$ is transformed to have zero mean and unit variance across all elements, resulting in $\mathbf{B}_{\rm std}$. Next, scaling and shifting are performed on $\mathbf{B}_{\rm std}$ along each column, which can be expressed as
$\mathbf{B}_{\rm pre}=\left(\bm{\alpha}\mathbf{1}_{K}^{\rm T}\right)*\mathbf{B}_{\rm std}+\bm{\beta}\mathbf{1}_{K}^{\rm T},$ (12)
where $\bm{\alpha}$ and $\bm{\beta}$ represent the trainable parameters for scaling and shifting, respectively, $\mathbf{1}_{K}$ is an all-ones vector of length $K$, and * denotes the Hadamard product of matrices. Finally, $\mathbf{B}_{\rm pre}$ is vectorized into $\mathbf{b}\in\mathbb{R}^{MK\times 1}$. The $i$th element of $\mathbf{b}$ is taken as the initial representation of the $i$th node, also denoted as $\mathbf{x}_{i}(0)$ in the context of the attention mechanism block.
2) Attention Mechanism Block: The attention mechanism allows the model to focus on specific parts of the input data by assigning varying importance to different components. To capture the interrelationship among nodes, a masked single-head attention mechanism is introduced, where masking enables the network to adapt to scenarios where the number of UEs varies. The attention mechanism block comprises $T=4$ hidden layers, with $\mathbf{x}_{i}(0)$ as the input to the first attention layer, the output sizes of $4$ layers are $32,64,64,64$, respectively. A detailed explanation of the attention mechanism block is provided in Subsection III-C.
3) Postprocessing Stage: Given the attention mechanism block’s output ${\mathbf{X}}(4)\in\mathbb{R}^{MK\times 64}$, whose $i$th row corresponds to $\mathbf{x}_{i}(4)$, a combination of linear and nonlinear processing is first applied. The resulting output can be described as
${\mathbf{Y}}_{1}={\mathcal{L}}_{2}({\rm ReLU}({\mathcal{L}}_{1}({\mathbf{X}}(4)))),$ (13)
where $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ represent two distinct linear transformations, each with input size of 64 and output size of 64, and ${\rm ReLU}(\cdot)$ denotes the rectified linear unit activation function.
Next, the node representation vectors undergo another linear transformation to map them into values. Following this, additional processing steps–similar to those in [7]–are applied, including the operations described in (14) and (15), giving
$\mathbf{y}_{2}=\exp(-{\rm SoftPlus}({\mathcal{L}}_{3}(\mathbf{Y}_{1})+6),$ (14)
where ${\rm SoftPlus}(\cdot)$ denotes the Softplus activation function, and ${\mathcal{L}}_{3}$ is a linear function that maps a vector to a scalar value. To extract the power control coefficients for the active users while ensuring that the results satisfy the constraints in (5), masking and projection operation ${\rm Proj}_{\mathcal{S}}(\cdot)$ in [13] are applied, yielding the power control coefficients as
$\hat{\mathbf{M}}={\rm Proj}_{\mathcal{S}}({\mathbf{Y}}_{2}\bm{\Phi}_{\rm mask}),$ (15)
where $\mathbf{Y}_{2}$ is obtained by reshaping the vector $\mathbf{y}_{2}$ into a matrix of size $M\times K$, and the diagonal matrix $\bm{\Phi}_{\rm mask}$ shares the same diagonal elements as $\bm{\Phi}$.
III-C Attention Mechanism for Graph Neural Network
Considering all layers of the attention mechanism share the same structure, differing only in the sizes of their input and output, the following description uses the $t$th layer as a representative example. Denoting the representation vector of the $i$th node as input in the $t$th layer as $\mathbf{x}_{i}(t-1)$, the output node representation vector is calculated by
$\mathbf{x}_{i}(t)=\mathrm{Norm}(\mathrm{ReLU}(\mathbf{y}_{{\rm AP},i}(t)+\mathbf{y}_{{\rm UE},i}(t))),i=1,\ldots,MK,$ (16)
where $\mathrm{Norm}(\cdot)$ denotes the normalization operation. Here $\mathbf{y}_{{\rm AP},i}(t)$ represents the component of the update through the attention mechanism, based on the information of nodes that are associated with the same APs as the $i$th node, and it can be expressed as
$\mathbf{y}_{{\rm AP},i}(t)=\mathcal{L}^{{\rm AP},t}_{1}\left(\mathbf{x}_{i}(t-1)\right)+$ (17)
$\mathcal{L}^{{\rm AP},t}_{5}\left(\sum\limits_{j\in\bm{\mathcal{V}}_{\rm AP},i}\psi_{ij}\alpha_{{\rm AP},ij}\left[\mathcal{L}^{{\rm AP},t}_{2}\left(\mathbf{x}_{j}(t-1)\right)+\mathcal{L}_{{\psi}}^{t}\left(\psi_{ij}\right)\right]\right),$
$\alpha_{{\rm AP},ij}=$ (18)
[…]
where $\mathcal{L}^{{\rm AP},t}_{q}$ ($q=1,\cdots,5$) denotes $5$ different linear transformations for capturing the information from AP-type edges in the $t$th layer, while $\mathcal{L}_{{\psi}}^{t}$ is a linear transformation used to capture pilot contamination information in the $t$th layer. Moreover, ${\bm{\mathcal{V}}}_{{\rm AP},i}$ denotes the set of nodes that share the same AP as the $i$th node, $\mathbf{a},\mathbf{b}\right>=\exp({\mathbf{a}^{\rm T}\mathbf{b}}/{\sqrt{D}})$, and $D$ is the length of vector $\mathbf{a}$.
Note that $\mathbf{y}_{{\rm UE},i}(t)$ is obtained through a similar process, but without considering pilot contamination. Specifically, $\mathcal{L}_{{\psi},t}\left(\cdot\right)$ is removed, as pilot contamination only arises between users.
IV Numerical Results
IV-A Simulation Setup
In this simulation, the large-scale fading coefficient $\beta_{mk}$ in $\rm dB$ is modeled as
$\beta_{mk}=\mathrm{PL}_{mk}+z_{mk},$ (19)
where $\mathrm{PL}_{mk}$ denotes the path loss in $\rm dB$, modeled using a three-slope model, and $z_{mk}$ accounts for shadow fading, following $\mathcal{CN}(0,\sigma_{\rm sh}^{2})$. The same large-scale fading parameters as in [2] are used to characterize the propagation path. The noise signal is simulated using the same parameters as in [7].
The lengths for the coherence block and pilot sequence are set to $T_{\rm c}=200$ and $T_{\rm p}=18$, respectively. A total of $T_{\rm p}$ orthogonal pilot sequences are allocated among UEs. The first $K$ sequences (corresponding to the actual number of UEs) are assigned, while for cases where $K>T_{\rm P}$, the remaining $K-T_{\rm p}$ sequences are randomly selected from the $T_{\rm P}$ available sequences. Simulation experiments are conducted in a CFmMIMO system with MMSE channel estimation in (3) and conjugate beamforming in (6).
IV-B Methods Compared and Performance Metrics
We compared the performance of the proposed method with two other power allocation methods: the APG method and the transformer-based method [^2] The proposed network and transformer-based network were trained on the NVIDIA H200 GPU.. The APG method, a traditional optimization-based method as described in [13], serves as a baseline to evaluate the performance of the proposed graph attention network, with the smoothness parameter $\lambda=3$ in (11). The transformer-based method of [7], a learning-based method also referred to as PAPC, is included for comparison as well. Considering that all existing GNN-based methods thus far are under the assumption of free pilot contamination, whereas our proposed method highlights its ability to handle pilot contamination, the GNN method is not included for comparison here.
The empirical cumulative distribution function (CDF) of the per-UE SE serves as a tool to visualize the SE distribution. In the context of the max-min fairness objective, a steeper CDF curve that shifts towards the right indicates a more desirable result, as it suggests improved fairness across UEs, as well as higher SEs for all UEs. Besides, we provide the running times of methods to assess their feasibility in real-time applications.
IV-C Performance Comparison in Different Scenarios
In this section, we compare the performance of $3$ methods across $5$ scenarios, demonstrating the adaptability and scalability of the proposed method. The parameter settings for $5$ scenarios are listed in Table I. For the first scenario, as an example, $K_{\rm max}=8$ UEs are served by $M=16$ APs, with both UEs and APs uniformly randomly distributed within a coverage area of $0.16$ ${\rm km}^{2}$. Since $K<T_{\rm p}$ in this case, pilot contamination does not occur. The number of UEs is not varying, thus $K=K_{\rm max}=K_{\rm min}$ for all sample data. If the number of UEs is varying, as in Scenarios 3 and 5, the number of UEs changes within $\left[K_{\rm min},K_{\rm max}\right]$. In Scenarios 2-5 with $K_{\rm max}>T_{\rm p}$, pilot contamination arises accordingly. All scenario results are tested on $500$ samples, and the number of training samples for each scenario is provided in Table I.
TABLE I: Parameters for $5$ Scenarios
Parameters Scen. 1 Scen. 2 Scen. 3 Scen. 4 Scen. 5
Coverage Area (${\rm km}^{2}$) 0.16 0.32 0.32 0.32 0.32
M 16 32 32 64 64
$K_{\rm max}$ 8 20 20 40 40
$K_{\rm min}$ 8 20 10 40 20
Pilot Contamination $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Varying UEs $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
Training Samples 50K 100K 800K 100K 800K
The CDF results for $5$ scenarios are presented in Fig. 2. Note that the proposed Graph ATtention network method is labeled as GAT in Figs. 2 and 3 for simplicity. It can be observed that the curve of the proposed method is to the right of those of both APG and PAPC in all scenarios, demonstrating its adaptability across different application scenarios, including variations in communication network size, pilot contamination, and whether the number of UEs is fixed or not. It is noteworthy that in Scenarios 3 and 5, the curves do not ascend as sharply as those of the other three scenarios. This is because the number of active UEs varies for each test sample, leading to a wider range of SE per UE.
To evaluate the flexibility of the proposed method, the trained models from Scenarios 3 and 5, which were trained with varying numbers of UEs, are tested on fixed UE settings, i.e., Scenario 3 with $K=20$ and Scenario 5 with $K=40$. The results, presented in Fig. 3, show that the models trained on datasets with varying numbers of UEs still demonstrate strong performance, even surpassing the APG method.
Table II presents the runtime for each method across different scenarios. The CPU measurements are obtained on a $64$-bit Windows system with $16$ GB RAM and an Intel Core i5 ($2.60$ GHz), while the GPU measurements are obtained using an NVIDIA A40. As shown, compared to $7.938\,[{\rm s}]$ processing time for APG in a scenario with $64$ APs and $40$ UEs, the proposed method achieves a processing time of under $46\,[{\rm ms}]$ on the GPU despite $552\,[{\rm ms}]$ on the CPU, highlighting the real-time performance of the learning-based method and demonstrating its potential for power control in real-time applications.
TABLE II: Runtime comparison between different methods (in seconds).
Scen. 1 Scen. 3 Scen. 5
APG on CPU $1.667$ $3.032$ $7.938$
Proposed on CPU $0.009$ $0.063$ $0.552$
Proposed on GPU $0.003$ $0.004$ $0.046$
Figure 2: Result comparison between $3$ methods across $5$ scenarios.
Figure 3: Performance comparison for the proposed method trained with a varying number of UEs but tested with fixed UEs.
V Conclusion
A graph attention method has been proposed for downlink power control in CFmMIMO systems. The proposed method not only effectively managed power control in the presence of pilot contamination but also demonstrated strong adaptability to varying communication network sizes and both fixed and dynamic numbers of UEs. Additionally, the neural network was trained in a self-supervised manner, eliminating the need for the labor-intensive process of labeling the large volume of samples. Experimental results demonstrated that the proposed method outperforms APG and transformer-based methods in terms of CDF for SE across various scenarios.